# HyperEvent: Learning Cohesive Events for Large-scale Dynamic Link Prediction

### v1

@@ B.1 Datasets
HyperEvent is evaluated using the Temporal Graph Benchmark (TGB), a comprehensive repository containing five temporal graph datasets for dynamic link prediction. Following TGB’s standardized protocol, all datasets are chronologically partitioned into training (70%), validation (15%), and test sets (15%) to ensure temporally meaningful evaluation. The experiments span five large-scale datasets representing diverse real-world dynamics: 1) Wiki (tgbl-wiki-v2): Captures bipartite user-page interactions from Wikipedia co-editing logs, featuring text-attributed edges. 2) Review (tgbl-review-v2): Contains weighted user-product interactions from Amazon electronics reviews (1997–2018). 3) Coin (tgbl-coin-v2): Documents cryptocurrency transactions during the 2022 Terra Luna market crash. 4) Comment (tgbl-comment): Models directed reply networks from Reddit (2005–2010), comprising over 44 million edges. 5) Flight (tgbl-flight-v2): Represents global airport traffic (2019–2022) with geo-attributed nodes and flight-number edges.
@@
As detailed in Table 3, these graphs exhibit substantial heterogeneity in scale (638K–18M nodes, 157K–67M edges), temporal dynamics (1,385–30M timestamps), and edge characteristics (weighted, directed, and attributed). The dynamic link prediction task consistently requires forecasting the next event given historical event streams.
All experiments implement the standardized evaluation protocols from TGB. This includes chronological train/validation/test splits and standardized negative queries for dynamic link prediction. While several datasets include node features or edge attributes/weights, such information is deliberately excluded. This choice isolates and highlights HyperEvent’s capacity to capture intrinsic event interdependencies, demonstrating the framework’s efficacy without relying on external features.
B.2 Evaluation Metric
The filtered Mean Reciprocal Rank (MRR) as specified by TGB is adopted, representing a ranking-based metric particularly suited for dynamic link prediction. Unlike Area Under the ROC Curve or Average Precision, which fail to adequately capture the relative ranking of positive edges against negatives, MRR directly measures the reciprocal rank of the true destination node among negative candidates. This aligns with recommendation systems and knowledge graph completion tasks where ranking quality is paramount. Crucially, all negative edges are predefined in TGB benchmark, ensuring fair comparison across methods.
Formally, for each positive query edge with prediction score $y_{\text{pos}}$, and a set of negative candidate scores $y_{\text{neg}}$, HyperEvent computes Optimistic rank: Number of negatives scoring strictly higher than $y_{\text{pos}}$. Pessimistic rank: Number of negatives scoring at least $y_{\text{pos}}$. The final rank $r$ is averaged:
@@ B.2 Evaluation Metric
Table 4: HyperEvent parameter configurations on different datasets
B.3 HyperEvent Parameters
@@
HyperEvent employs three key hyperparameters: $n_{\text{neighbor}}$, $n_{\text{latest}}$, and $n_{\text{segment}}$. The hyperparameter $n_{\text{segment}}$ primarily governs computational efficiency by segmenting the event stream for parallel processing. Selection of $n_{\text{segment}}$ balances GPU memory constraints with training duration relative to dataset scale. During evaluation, which involves processing multiple negative samples concurrently, a smaller $n_{\text{segment}}$ value is typically used compared to training. For the Wiki, Review, and Comment datasets, the optimal $n_{\text{neighbor}}$ was searched within {10, 15, 20}, with $n_{\text{latest}}$ fixed at 10. Observations indicated a preference for larger neighborhoods on Coin and Flight datasets, prompting a search over {10, 15, 20, 30, 40, 50} for $n_{\text{neighbor}}$. The hyperparameter $n_{\text{latest}}$ was optimized from {5, 8, 10} when fixing $n_{\text{neighbor}}=10$. Detailed hyperparameter configurations for all datasets are provided in Table 4.
B.4 Transformer Architecture
@@
The hyper-event discriminator utilizes a Transformer module. As this component does not represent a core innovation, a fixed architectural configuration was employed across all datasets without dataset-specific tuning. The settings include: a maximum sequence length of 512 for position encoding, position encoding dimension of 64, Transformer hidden dimension of 64, 3 layers, 4 attention heads, dropout rate of 0.1, and GELU activation functions.
B.5 Training Protocol
@@
Optimization employed Adam with a fixed learning rate of 0.0001. The model was trained for 50 epochs on the Wiki dataset and 5 epochs on all other datasets. Model selection was based on performance on the validation set, with final results reported on the test set using the MRR metric. The mean and standard deviation of MRR over 5 independent runs with different random seeds are reported to ensure statistical reliability.
B.6 Hardware
@@
All experiments were conducted on a high-performance computing system equipped with an Intel Core i7-14700KF CPU, 64 GB RAM, and an NVIDIA GeForce RTX 4090 GPU with 24 GB VRAM.
B.7 Ablation Study
@@
Influence of $n_{\text{latest}}$
@@
This behavior can be attributed to the definition of $n_{\text{latest}}$, which directly governs the number of recent events aggregated into a hyper-event. Datasets primarily rely on $n_{\text{neighbor}}$ to encode long- and short-term historical preferences; $n_{\text{neighbor}}$ effectively captures node-level temporal dependencies, thereby reducing the influence of $n_{\text{latest}}$ on overall predictive performance. Furthermore, excessively increasing $n_{\text{latest}}$ escalates computational overhead due to the linear scaling of event aggregation in hyper-event formation, without delivering commensurate gains in accuracy for most scenarios. Consequently, adopting $n_{\text{latest}}=10$ strikes an optimal balance, providing robust predictive capabilities while maintaining computational efficiency across diverse dynamic graph environments.
@@
Table 5: MRR of HyperEvent with/without Enhanced Correlation Vector
Without the Enhanced Correlation Vector
@@
This ablation evaluates the Enhanced Correlation Vector’s contribution by comparing model performance when replacing the full 12-dimensional correlation vector with a reduced version using only 1-hop neighborhood overlap distance. As Table 5 indicates, removing multi-hop structural correlations causes consistent and substantial performance degradation. The MRR drops range from 4.06% on Wiki to a drastic 67.91% on Review, with all five datasets exhibiting declines exceeding 10% except Wiki. This degradation proves particularly severe for datasets like Review and Comment, where node interactions involve complex higher-order dependencies.
@@
Two key factors explain this sensitivity. First, limiting correlation modeling to immediate (1-hop) neighborhoods overlooks critical structural contexts captured by 0-hop (node identity correlation) and 2-hop (extended topology correlation) features. For example, on Review, where users interact sequentially across diverse products, 0-hop identity matching helps confirm recurring user/item relationships, while 2-hop signals reveal indirect influence chains. Second, the impoverished 4-dimensional vector lacks granularity to discriminate between genuine hyper-events and coincidental event aggregations. Consequently, the model frequently misclassifies valid query events that rely on multi-scale evidence, especially for inductive cases involving unseen nodes where 1-hop patterns alone are insufficient to confirm hyper-event integrity. These results validate that comprehensive correlation modeling—spanning local to global neighborhood contexts—is indispensable for robust hyper-event recognition.
Without the Efficient Training Algorithm
@@
Table 6 quantitatively compares per-epoch training durations for HyperEvent with and without the proposed efficient training algorithm across five temporal graph datasets. A striking observation emerges: enabling this algorithm consistently reduces training times by over 95% across all datasets, achieving relative savings exceeding 98% on four benchmarks. Notably, the degree of acceleration correlates with the segmentation parameter $n_{\text{segment}}$. Higher $n_{\text{segment}}$ values yield more significant efficiency gains, as evidenced by the 98.07% saving on Flight ($n_{\text{segment}}=50$) versus 95.03% on Comment ($n_{\text{segment}}=20$).
@@
This dramatic improvement stems from the algorithm’s core design: partitioning the monolithic event stream into $n_{\text{segment}}$ independent segments linearly reduces the effective event sequence length processed within each iteration from $|\mathcal{E}|$ to $\frac{|\mathcal{E}|}{n_{\text{segment}}}$. Simultaneously, tensorized implementation can fully leverage the massive parallel computing capabilities of GPUs while avoiding multithreading overhead. Consequently, the computational complexity per epoch scales inversely with $n_{\text{segment}}$, explaining the near-linear time reduction observed. Such efficiency is imperative for scaling HyperEvent to massive real-world temporal graphs, where raw training times without segmentation proved prohibitive (e.g., exceeding 55 hours per epoch on Flight). This approach transforms HyperEvent from a conceptual framework into a practically deployable solution for industrial-scale dynamic graphs.
@@
Table 6: Per-epoch Training Time of HyperEvent with/without Efficient Training Algorithm

### v2

@@ IV-A Datasets
HyperEvent is evaluated using the Temporal Graph Benchmark (TGB) [47], a comprehensive repository containing five temporal graph datasets for dynamic link prediction. Following TGB’s standardized protocol, all datasets are chronologically partitioned into training (70%), validation (15%), and test sets (15%) to ensure temporally meaningful evaluation. The experiments span five large-scale datasets representing diverse real-world dynamics: 1) Wiki (tgbl-wiki-v2): Captures bipartite user-page interactions from Wikipedia co-editing logs, featuring text-attributed edges[10]. 2) Review (tgbl-review-v2): Contains weighted user-product interactions from Amazon electronics reviews (1997–2018)[48]. 3) Coin (tgbl-coin-v2): Documents cryptocurrency transactions during the 2022 Terra Luna market crash[49]. 4) Comment (tgbl-comment): Models directed reply networks from Reddit (2005–2010), comprising over 44 million edges[50]. 5) Flight (tgbl-flight-v2): Represents global airport traffic (2019–2022) with geo-attributed nodes and flight-number edges[51].
@@
As detailed in Table I, these graphs exhibit substantial heterogeneity in scale (638K–18M nodes, 157K–67M edges), temporal dynamics (1,385–30M timestamps), and edge characteristics (weighted, directed, and attributed). The dynamic link prediction task consistently requires forecasting the next event given historical event streams.
IV-B Compared Methods
HyperEvent is compared against the following temporal graph representation learning and dynamic link prediction methods, as introduced in Section I:
@@ IV-B Compared Methods
Deep learning methods: JODIE [10], DyRep [11], TGAT [43], TGN [9], CAWN [15], NAT [14], TCL [46], GraphMixer [17], DyGFormer [16], TNCN [12], CTAN [13], TPNet [45].
IV-C Evaluation Metric
The filtered Mean Reciprocal Rank (MRR) as specified by TGB [47] is adopted, representing a ranking-based metric particularly suited for dynamic link prediction. Unlike Area Under the ROC Curve or Average Precision, which fail to adequately capture the relative ranking of positive edges against negatives, MRR directly measures the reciprocal rank of the true destination node among negative candidates. This aligns with recommendation systems and knowledge graph completion tasks where ranking quality is paramount. Crucially, all negative edges are predefined in TGB benchmark, ensuring fair comparison across methods.
Formally, for each positive query edge with prediction score $y_{\text{pos}}$, and a set of negative candidate scores $y_{\text{neg}}$, HyperEvent computes Optimistic rank: Number of negatives scoring strictly higher than $y_{\text{pos}}$. Pessimistic rank: Number of negatives scoring at least $y_{\text{pos}}$.
The final rank $r$ is averaged:
$r=\frac{1}{2}(\text{optimistic\_rank}+\text{pessimistic\_rank})+1.$ (18)
@@ IV-C Evaluation Metric
IV-D Experimental Setup
TABLE II: HyperEvent Parameter Configurations on Different Datasets
@@ IV-D Experimental Setup
Dataset Configuration. All experiments implement the standardized evaluation protocols from TGB [47]. This includes chronological train/validation/test splits and standardized negative queries for dynamic link prediction. While several datasets include node features or edge attributes/weights, such information is deliberately excluded. This choice isolates and highlights HyperEvent’s capacity to capture intrinsic event interdependencies, demonstrating the framework’s efficacy without relying on external features.
@@
HyperEvent Parameters. HyperEvent employs three key hyperparameters: $n_{\text{neighbor}}$, $n_{\text{latest}}$, and $n_{\text{segment}}$. The hyperparameter $n_{\text{segment}}$ primarily governs computational efficiency by segmenting the event stream for parallel processing. Selection of $n_{\text{segment}}$ balances GPU memory constraints with training duration relative to dataset scale. During evaluation, which involves processing multiple negative samples concurrently, a smaller $n_{\text{segment}}$ value is typically used compared to training. For the Wiki, Review, and Comment datasets, the optimal $n_{\text{neighbor}}$ was searched within {10, 15, 20}, with $n_{\text{latest}}$ fixed at 10. Observations indicated a preference for larger neighborhoods on Coin and Flight datasets, prompting a search over {10, 15, 20, 30, 40, 50} for $n_{\text{neighbor}}$. The hyperparameter $n_{\text{latest}}$ was optimized from {5, 8, 10} when fixing $n_{\text{neighbor}}=10$. Detailed hyperparameter configurations for all datasets are provided in Table II.
@@
Transformer Architecture. The hyper-event discriminator, implemented via Eq. (16), utilizes a Transformer module. As this component does not represent a core innovation, a fixed architectural configuration was employed across all datasets without dataset-specific tuning. The settings include: a maximum sequence length of 512 for position encoding, position encoding dimension of 64, Transformer hidden dimension of 64, 3 layers, 4 attention heads, dropout rate of 0.1, and GELU activation functions.
@@
Training Protocol. Optimization employed Adam with a fixed learning rate of 0.0001. The model was trained for 50 epochs on the Wiki dataset and 5 epochs on all other datasets. Model selection was based on performance on the validation set, with final results reported on the test set using the MRR metric. The mean and standard deviation of MRR over 5 independent runs with different random seeds are reported to ensure statistical reliability.
@@
Hardware. All experiments were conducted on a high-performance computing system equipped with an Intel Core i7-14700KF CPU, 64 GB RAM, and an NVIDIA GeForce RTX 4090 GPU with 24 GB VRAM.
MRR($\mathbf{\times 100}$). For clearer visualization in tables and figures, all presented MRR values are scaled by a factor of 100; textual references to MRR subsequently implicitly denote these scaled values.
TABLE III: MRR of methods on different datasets
Model Wiki Review Coin Comment Flight
EdgeBank 49.5 2.3 45.2 12.9 16.7
[…]
IV-E Methods Comparison
Table III presents the MRR of HyperEvent against 13 methods. HyperEvent demonstrates strong performance, achieving competitive results on four datasets: Wiki (80.9), Coin (77.3), Comment (75.9), and Flight (87.7), where it is outperformed only by the TPNet model. This performance is achieved through a simple framework that leverages relative structural encoding features and a lightweight transformer classifier. The fact that this straightforward approach achieves competitive results highlights the effectiveness of relative structural encoding in dynamic link prediction.
1) High Temporal Surprise: With a surprise metric of 0.987 (the highest among datasets), Review exhibits highly unpredictable interactions. HyperEvent’s reliance on historical event patterns becomes less effective when future events deviate substantially from past structures.
2) Sparse Event Context: Predicting ”which product a user reviews next” relies heavily on individual user behavior rather than inter-event dependencies. This misalignment with HyperEvent’s strength in modeling group dynamics reduces its relative advantage.
This result is instructive for future research: it clearly delineates a scenario (high surprise, sparse context) where a simple, relative-structural-based baseline may be insufficient, and more specialized models are warranted. Thus, HyperEvent serves not only as a performance reference point but also as a diagnostic tool to identify problem characteristics that demand more sophisticated solutions.
The strong performance on Coin, Comment, and Flight—datasets featuring dense event streams and strong temporal dependencies—reinforces HyperEvent’s effectiveness in scenarios where event interdependencies drive evolutionary patterns. HyperEvent establishes a new, simple, and effective baseline for temporal graph representation learning.
IV-F Effectiveness of the Training Algorithm
TABLE IV: Efficiency of HyperEvent on Flight Dataset
Method Metric
Training Time
per Epoch
GPU Memory MRR
TNCN 10h 34m 37s 15,189MB 82.0
HyperEvent(ours) 1h 4m 33s 2,346MB 87.7
Improvement 89.83% $\downarrow$ 84.55% $\downarrow$ 6.95% $\uparrow$
Figure 5: GPU memory usage (MB, right axis) and per-epoch training duration (seconds, left axis) of HyperEvent across varying segment counts ($n_{\text{segment}}$).
The efficiency gains of HyperEvent are quantified in Table IV, contrasting HyperEvent with TNCN. HyperEvent achieves a remarkable 89.83% reduction in per-epoch training time (1h 4m 33s vs. TNCN’s 10h 34m 37s) and an 84.55% reduction in GPU memory consumption (2,346MB vs. 15,189MB), while simultaneously improving prediction accuracy with a 6.95% higher MRR (87.7 vs. 82.0). These results underscore the practicality of HyperEvent as a baseline. Its efficiency is not merely an ancillary benefit but a core feature, ensuring that it can be rapidly trained and evaluated on large-scale graphs, making it an accessible and scalable reference point for the research community.
The efficiency stems from three key aspects: First, segmenting the event stream into parallelizable segments reduces the iterations per epoch by a factor of $n_{\text{segment}}$. Second, a tensorized implementation leverages GPU parallelism directly, bypassing multi-threading bottlenecks and CPU-GPU communication latency. Third, the low-dimensional design of the relative structural encoding vectors (only 12 dimensions) and the lightweight classification model significantly reduce the memory footprint.
Collectively, these results demonstrate that HyperEvent establishes a new benchmark for computationally efficient and effective temporal graph learning. It provides a compelling baseline that is simple in its core idea and highly efficient in practice, setting a clear reference point for future work.
IV-G Modulating Long- and Short-Term Preferences via $n_{\text{neighbor}}$ and $n_{\text{latest}}$
IV-G1 Influence of $n_{\text{neighbor}}$
This divergent behavior originates from the role of $n_{\text{neighbor}}$ in defining the temporal scope for constructing event sequences, as introduced in Section III-C1. Here, $n_{\text{neighbor}}$ represents the number of most recent interaction partners stored chronologically in the adjacency table, which directly determines the historical length considered for each node. In HyperEvent, this parameter modulates the input to the relative structural encoding calculations from equations (9) to (LABEL:eq_correlation), thereby balancing long-term preferences, such as persistent node affinities over extended timelines, against short-term preferences that focus on immediate, transient interactions. Datasets like Coin and Flight, characterized by slowly evolving patterns, achieve higher MRR with larger $n_{\text{neighbor}}$ values, as they leverage richer historical context to identify event patterns. In contrast, datasets such as Review and Comment, which involve rapid, high-frequency interactions, perform better with shorter histories, as excessive context may introduce noise that obscures recent relevant events. The findings underscore the adaptability of HyperEvent in capturing dynamic graph dynamics, highlighting how tuning $n_{\text{neighbor}}$ effectively addresses the inherent temporal heterogeneities across applications.
@@ IV-G1 Influence of $n_{\text{neighbor}}$
IV-G2 Influence of $n_{\text{latest}}$
@@
This behavior can be attributed to the definition of $n_{\text{latest}}$, which directly governs the number of recent events aggregated into a sequence. As elaborated in Section IV-G1, datasets primarily rely on $n_{\text{neighbor}}$ to encode long- and short-term historical preferences; $n_{\text{neighbor}}$ effectively captures node-level temporal dependencies, thereby reducing the influence of $n_{\text{latest}}$ on overall predictive performance. Furthermore, excessively increasing $n_{\text{latest}}$ escalates computational overhead due to the linear scaling of event aggregation in sequence formation, without delivering commensurate gains in accuracy for most scenarios. Consequently, adopting $n_{\text{latest}}=10$ strikes an optimal balance, providing robust predictive capabilities while maintaining computational efficiency across diverse dynamic graph environments.
IV-H Ablation Study
TABLE V: MRR of HyperEvent with/without Enhanced Correlation Vector
@@ IV-H Ablation Study
IV-H1 Without the Relative Structural Encoding Vector
@@
This ablation evaluates the contribution of the relative structural encoding vector by comparing model performance when replacing the full 12-dimensional relative structural encoding vector (integrating equations 13, 12, and 14) with a reduced version using only 1-hop neighborhood overlap distance (Equation 13). As Table V indicates, removing multi-hop structural features causes consistent and substantial performance degradation. The MRR drops range from 4.06% on Wiki to a drastic 67.91% on Review, with all five datasets exhibiting declines exceeding 10% except Wiki. This degradation proves particularly severe for datasets like Review and Comment, where node interactions involve complex higher-order dependencies.
@@
Two key factors explain this sensitivity. First, limiting encoding to immediate (1-hop) neighborhoods overlooks critical structural contexts captured by 0-hop (node identity) and 2-hop (extended topology) features. For example, on Review, where users interact sequentially across diverse products, 0-hop identity matching helps confirm recurring user/item relationships, while 2-hop signals reveal indirect influence chains. Second, the impoverished 4-dimensional vector lacks granularity to discriminate between genuine event sequences and coincidental event aggregations. Consequently, the model frequently misclassifies valid query events that rely on multi-scale evidence, especially for inductive cases involving unseen nodes where 1-hop patterns alone are insufficient to confirm sequence integrity. These results validate that comprehensive relative structural encoding—spanning local to global neighborhood contexts—is beneficial for effective event sequence recognition.
IV-H2 Without the Efficient Training Algorithm
@@
Table VI quantitatively compares per-epoch training durations for HyperEvent with and without the proposed efficient training algorithm across five temporal graph datasets. A striking observation emerges: enabling this algorithm consistently reduces training times by over 95% across all datasets, achieving relative savings exceeding 98% on four benchmarks. Notably, the degree of acceleration correlates with the segmentation parameter $n_{\text{segment}}$. Higher $n_{\text{segment}}$ values yield more significant efficiency gains, as evidenced by the 98.07% saving on Flight ($n_{\text{segment}}=50$) versus 95.03% on Comment ($n_{\text{segment}}=20$).
@@
This dramatic improvement stems from the algorithm’s core design: partitioning the monolithic event stream into $n_{\text{segment}}$ independent segments linearly reduces the effective event sequence length processed within each iteration from $|\mathcal{E}|$ to $\frac{|\mathcal{E}|}{n_{\text{segment}}}$. Simultaneously, tensorized implementation in Section III-D can fully leverage the massive parallel computing capabilities of GPUs while avoiding multithreading overhead. Consequently, the computational complexity per epoch scales inversely with $n_{\text{segment}}$, explaining the near-linear time reduction observed. Such efficiency is imperative for scaling HyperEvent to massive real-world temporal graphs, where raw training times without segmentation proved prohibitive (e.g., exceeding 55 hours per epoch on Flight). This approach transforms HyperEvent from a conceptual framework into a practically deployable solution for industrial-scale dynamic graphs.
TABLE VI: Per-epoch Training Time of HyperEvent with/without Efficient Training Algorithm
@@
V conclusion
In this work, we observe that many recent advances in temporal graph representation learning for dynamic link prediction are evaluated without comparison to a consistently strong and efficiently reproducible baseline. This makes it challenging to assess whether performance gains stem from genuinely useful innovations or from incidental architectural complexity. To help address this, we introduce HyperEvent, a simple and lightweight baseline that encodes relative structural patterns among recent events using handcrafted features and a minimal transformer backbone.
Rather than attempting to model complex temporal dynamics or causal relationships, HyperEvent focuses on capturing basic local structural regularities in the evolving graph through relative positional and relational encodings. Surprisingly, this straightforward approach achieves competitive or superior performance across five established TGB benchmarks, often with significantly fewer parameters and faster training than more elaborate learned models.
We do not claim that HyperEvent captures deep temporal semantics or event level dependencies. Its effectiveness suggests that simple structural priors, when carefully designed, can serve as a surprisingly strong foundation. We hope HyperEvent can act as a pragmatic reference point for future work, helping to calibrate progress, reduce redundant complexity, and encourage clearer ablation studies.
Limitations remain, particularly in settings where graph evolution is highly irregular or lacks discernible local structure. We look forward to seeing future methods that reliably surpass this baseline under such conditions. Ultimately, our goal is not to propose a final solution, but to offer a humble and reproducible starting point that helps the community build more meaningfully on solid ground.
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.